### v2

@@ 4.1 Data
We first consider a univariate design simulated using data from Lee (2008), as the Lee (2008) data is a popular choice to conduct simulations in the RD literature. Lee (2008) studies the effect of being the current incumbent party in a district on the votes obtained in the district’s next election. The running variable is univariate and measures the difference in vote share between the Democratic candidate and their strongest opponent in the current election, and the outcome variable is the Democratic vote share in the next election. This univariate design normalizes the assignment cutoff to 00, and units with $X_{i}\geq 0$ are treated (i.e., winning the current election).
@@
We then consider a bivariate design simulated using data from Keele and Titiunik (2015), who study the effect of television advertising on election turnout, where the running variable is the two-dimensional geographic coordinate of a voter’s location and the outcome of interest is election turnout and takes binary values. Treatment status is determined by a geographic media market boundary. Their data also contains several other outcome variables for placebo analysis. In addition to election turnout, we include age and housing price per square foot in our simulation to cover both discrete and continuous outcome variables. Keele and Titiunik (2015) consider three equally spaced points along the treatment boundary, depicted as black squares in Figure 3, to be the points at which they estimate the treatment effect Without loss of generality, we pick the middle point, circled in green in Figure 3, to be the boundary point of interest in our simulation.
@@
We consider designs up to two scores in our Monte Carlo study because the current R implementation of the minimax-optimal estimator $\widehat{\tau}_{\texttt{mm}}(x^{c};\mathcal{B})$ in (7) provided by Imbens and Wager (2019) only accepts up to two scores. But in work in progress, we illustrate how to use the two forest-based estimators in RD designs with more than two scores to explore heterogeneous treatment effects along the treatment boundary in an empirical application.
4.2 Generating Simulation Data
To evaluate the aforementioned nonparametric estimators, we design a systematic Monte Carlo study with data generating processes built both from functional forms that we specify, and, following the proposal of Athey et al. (2021), from Wasserstein Generative Adversarial Networks (WGANs) with a high degree of complexity that can closely mimic the observed data.
@@ 4.2 Generating Simulation Data
We first consider fitting the observed data with a polynomial, which is a popular choice in the RD literature to evaluate estimators. Imbens and Kalyanaraman (2012), Calonico et al. (2014), and Calonico et al. (2019) all consider simulation data generated by using the Lee (2008) data to fit two $5$-th order global polynomials separately for the treated and control populations, which we also consider in our simulation. For the Keele and Titiunik (2015) data, we fit a $3$-rd order polynomial for continuous outcomes (i.e., age and housing price) and a logistic regression using, again, a $3$-rd order polynomial inside the logit link function for the binary outcome turnout, separately for the treated and control populations. [^7] We also tried higher order polynomials and the results do not qualitatively differ. Details on these polynomial DGPs can be found in Appendix B.1.
@@
As argued by Athey et al. (2021), Monte Carlo studies are more credible when the data used to evaluate the performance of econometric methods does not come from researcher-chosen functional forms. Following their proposal, we also consider simulation data generated from WGAN, which trains two neural networks—one called the generator and the other called the critic—in an adversarial setting to minimize the Wasserstein distance between a real dataset and the artificial dataset generated from the neural nets. We refer the reader to Athey et al. (2021) for a technical introduction to WGAN.
@@
Using WGAN to generate data to evaluate estimators for treatment effects, however, is problem-specific, since we never directly observe the treatment effect in the data, but rather we “identify” it under a set of identification assumptions. Athey et al. (2021) provide an example of how to generate simulation data to evaluate estimators of treatment effects under unconfoundedness, where they train two conditional WGANs to learn $X_{i}|D_{i}$ and $Y_{i}|(X_{i},D_{i})$ from the real sample. They then draw a fake sample of treatment status and covariates from the WGAN-generated distribution of $X_{i}|D_{i}$, and for each $(X_{i},D_{i})$ drawn, an outcome is drawn from the WGAN-generated distribution of $Y_{i}|(X_{i},D_{i})$. Under unconfoundedness, $Y_{i}|(X_{i},D_{i}=d)$ identifies $Y_{i}(d)|X_{i}$ and hence for each $Y_{i}$ drawn, its counterfactual outcome is generated by mutating the corresponding treatment status $D_{i}$ to $(1-D_{i})$ but keeping the same $X_{i}$, i.e., the counterfactual outcome is drawn from the WGAN-generated distribution of $Y_{i}|(X_{i},1-D_{i})$. The true average treatment effect is simulated by averaging the difference between the pair of generated outcome and its counterfactual over a large WGAN-generated sample.
@@
Our setting, with a different design and identification assumption for the treatment effect of interest, requires a different way of using WGAN to generate suitable simulation data. Specifically, the treatment effects are only identified for $X_{i}\in\mathscr{B}$, so we need to simulate the treatment effect at each treatment boundary point of interest, instead of simply taking averages as in the unconfoundedness setting. To do so, we first train two conditional WGANs to learn $X_{i}|D_{i}$ and $Y_{i}|(X_{i},D_{i})$ from the real sample, as in Athey et al. (2021). But for a given treatment boundary point $x^{c}\in\mathscr{B}$, we draw a separate artificial sample to simulate the true treatment effect at $x^{c}$: after training the WGAN to learn the distribution of $Y_{i}|(X_{i},D_{i})$, we generate 100,000 outcomes from $Y_{i}|(X_{i}=x^{c},D_{i}=1)$ and another 100,000 outcomes from $Y_{i}|(X_{i}=x^{c},D_{i}=0)$, and take the average of their differences as the true treatment effect at $x^{c}$. [^8] We note that WGANs do not directly yield closed-form estimates of the fitted distribution, but rather generate a new, artificial sample of arbitrary size that closely mimics the real sample. This is why we need to simulate the true treatment effect at a given treatment boundary point $x^{c}\in\mathscr{B}$, unlike in the polynomial case where we can directly calculate the true treatment effects from the coefficients of the fitted polynomials evaluated at $x^{c}$. Figures 2 and 3 visualize the actual data and the simulated data from the polynomial specification and WGAN. In the Appendix, we include additional visualization of the WGAN-generated DGPs in Figures B.1B.7, and we report in Table B.2 the Wasserstein distance between the real data and the WGAN data.
@@
Figure 2: Lee (2008): Democratic victory margin in the current election against the vote share in the next election. Panel A shows the observed data. Panel B plots the model data fitted using a 5-th order polynomial with different coefficients for the treated and control observations. Panel C plots the WGAN data. The cutoff is normalized to 0. The figure is generated using the rdrobust package (Calonico et al. (2014)) with default smoothing.
We use the default options available from the wgan package of Athey et al. (2021) to train the WGANs; see Appendix B.3 for details. Following Section 5 of Athey et al. (2021), we conduct a thorough robustness check to assess the variability of the Monte Carlo results with respect to sampling variation and WGAN tuning parameters, with robustness results reported in Appendix B.5.
@@
Figure 3: Visualization of the sample data in Keele and Titiunik (2015) (left) as well as simulated polynomial and WGAN DGPs (middle and right, respectively). The outcome variable is the binary election turnout. Coordinates below the black curve (treatment boundary) are treated, and the green circle shows the boundary point at which we estimate the treatment effect. To visualize the polynomial DGPs, 2,500 coordinates are sampled with replacement from the original sample plus a noise term $N(0,0.01^{2})$.
@@ 4.3 Monte Carlo Results
We report the Monte Carlo simulation results in Table 1. All simulations are implemented in R, where we use the rdrobust package of Calonico et al. (2014) for local linear regressions, the optrdd package of Imbens and Wager (2019) for the minimax-optimal estimator, and the grf package of Athey et al. (2019) for honest regression forests and local linear forests—all with default options without tuning, except we specify a few hyperparameters for the two forest-based estimators for reasons we explain in Appendix B.6.
@@
Table 1: Monte Carlo simulation results. For each method, we vary the sample size $n\in\{1000,5000,10000,50000,100000\}$ and report the average bias, variance, and $95\%$ coverage rate across 1,000 Monte Carlo replications. Each panel corresponds to a different outcome variable in Section 4.1, with the true treatment effect in parentheses either calculated exactly (from the polynomial fit) or simulated as described in Section 4.2 (for the WGAN DGPs).
@@
Local linear regressions perform well in univariate settings, although they slightly undercover in the polynomial DGP (Panel A, Columns 1-5 of Table 1). [^9] This undercoverage is consistent with the simulation result ($90.6\%$ coverage) of Calonico et al. (2019), where they use the same polynomial DGP but with $n=1000$ for a total of $5000$ replications; see Table SA-1, Row 2 and Column 3 in their supplemental appendix. We have done our simulation using their STATA code and obtained similar results to those reported in Table 1. However, they are in general noisier in bivariate settings, where we follow Keele and Titiunik (2015) to first collapse the bivariate score using rescaled Euclidean (chordal) distance and then run rdrobust, especially in small samples. Although the variance of local linear regressions shrinks as the sample size gets large, their bias does not shrink fast enough relative to the diminishing variance in those bivariate WGAN DGPs (Panels F-H, Columns 1-5 of Table 1), and therefore the coverage rate drops as the sample size increases. This problem arises possibly due to the zero-density issue of the collapsed multivariate score at the new univariate cutoff, as we point out in Section 3.1.1 and empirically illustrate in the last two histograms of Figure 1, cautioning against univariate transformations of multivariate scores.
@@
The minimax-optimal estimator of Imbens and Wager (2019) has a relatively low bias but tends to overcover, which is expected since it is designed for minimizing the worst-case MSE and the construction of confidence intervals aims at achieving uniform coverage over a class of CEFs explained in Section 3.2. Nevertheless, its performance depends crucially on accurate estimates of the second derivative bound $\mathcal{B}$: for the two artificial datasets built using the Lee (2008) data (Panels A-B of Table 1), the minimax-optimal estimator undercovers severely, especially for the polynomial DGP in Panel A. There, we follow the recommendation of Imbens and Wager (2019) to estimate the CEF globally using a second-order polynomial and multiply the maximal estimated curvature by $4$. But this turns out to yield a too small estimate of $\mathcal{B}$; Table B.1 in Appendix B.2 shows how the simulation results change with the constant we multiply, and it is only after multiplying the maximal estimated curvature by $30$ do we get an estimate that is close to the true (least upper) bound $\mathcal{B}$. Of course, one could argue that the global second-order polynomial is a bad model to start with; in the bivariate design, we follow their working paper (Imbens and Wager, 2018, p.21) and estimate $\mathcal{B}$ using the $95$-th percentile (over all sample points) of the operator norm curvature via a cross-validated ridge regression with interacted $7$-th order natural splines, which yields good coverage results (Panels C-H, Columns 6-10 of Table B.1). However, without knowledge about the true DGP, choosing a “good” model is not a straightforward problem, and estimating second derivatives is more difficult when the dimension of the score increases. In addition, the current implementation of optrdd accepts up to two scores only, limiting its application more broadly.
@@
Honest regression forests and local linear forests have comparable performance to local linear regressions in the univariate design (Panels A-B, Columns 11-20 of Table 1), though their variances are slightly higher. On the other hand, regression forests are in general more biased than local linear forests, especially in small samples, which is a known boundary problem for regression forests. In addition, the variance of regression forests tends to be smaller than that of local linear forests, exacerbating the undercoverage problem. As we discussed in Section 3.3.2, local linear forests are better at capturing smooth signals via a linear correction, and its advantage over regression forests is especially pronounced when the underlying DGP exhibits strong linear trends near the boundary point at which we estimate the treatment effect. In Appendix B.5, we examine the robustness of the results in Table 1 to the number of training epochs, which is a tuning parameter of WGAN, and increasing the number of training epochs leads to stronger linear trends near the boundary point (see Figure B.7); in these cases, the performance of regression forests deteriorates, whereas local linear regressions have robust performance. With that said, these two forest-based methods are valid alternatives to existing methods for multivariate RD designs that can flexibly model any given number of scores, and empirical researchers may leverage and tailor them accordingly to fit different empirical settings.

### v3

@@ 4.1 Data
We first consider a univariate design simulated using data from Lee (2008), which is a popular choice in the RD literature to conduct simulations. Lee (2008) studies the effect of being the current incumbent party in a district on the votes obtained in the district’s next election. The running variable is univariate and measures the difference in vote share between the Democratic candidate and their strongest opponent in the current election, and the outcome variable is the Democratic vote share in the next election. This univariate design normalizes the assignment cutoff to 00, and units with $X_{i}\geq 0$ are treated (i.e., winning the current election).
@@
We then consider a bivariate design simulated using data from Keele and Titiunik (2015), who study the effect of television advertising on election turnout, where the running variable is the two-dimensional geographic coordinate of a voter’s location and the outcome of interest is election turnout and takes binary values. Treatment status is determined by a geographic media market boundary. Their data also contains several other outcome variables for placebo analysis. In addition to election turnout, we include age and housing price per square foot in our simulation to cover both discrete and continuous outcome variables. Keele and Titiunik (2015) consider three equally spaced points along the treatment boundary, depicted as black squares in Figure 3, to be the points at which they estimate the treatment effect. For simplicity, we pick the middle point, circled in green in Figure 3, to be the boundary point of interest in our simulation.
@@
We consider designs up to two scores in our Monte Carlo study because the current R implementation of the minimax-optimal estimator $\widehat{\tau}_{\texttt{mm}}(x^{c};\mathcal{B})$ in (6) provided by Imbens and Wager (2019) only accepts up to two scores. In an empirical application in Section 5, we illustrate how to use the forest-based estimators in an RD design with more than two scores to explore heterogeneous treatment effects along the treatment boundary.
4.1.1 Generating Simulation Data
@@
We consider data generating processes built both from functional forms that we specify, and, following the proposal of Athey et al. (2021), from Wasserstein Generative Adversarial Networks (WGANs) with a high degree of complexity that can closely mimic the observed data. We first fit the observed data with a polynomial: Imbens and Kalyanaraman (2012), Calonico et al. (2014), and Calonico et al. (2019) all use the Lee (2008) data to generate simulation data by fitting two $5$-th order global polynomials separately for the treated and control populations, which we also use in our simulation. For the Keele and Titiunik (2015) data, we fit a $3$-rd order interacted polynomial for continuous outcomes (i.e., age and housing price) and a logistic regression using a $3$-rd order interacted polynomial inside the logit link function for the binary outcome turnout, separately for the treated and control populations. Details on these polynomial DGPs can be found in Appendix B.1.
@@
As argued by Athey et al. (2021), Monte Carlo studies are more credible when the data used to evaluate the performance of econometric methods does not come from researcher-chosen functional forms. Following their proposal, we also consider simulation data generated from WGAN, which trains two neural networks—one called the generator and the other called the critic—in an adversarial setting to minimize the Wasserstein distance between a real dataset and the artificial dataset generated from the neural nets. We refer the reader to Athey et al. (2021) for an introduction to WGAN.
@@
Using WGAN to generate data to evaluate estimators for treatment effects, however, is task-specific, since we never directly observe the treatment effect in the data, but rather we “identify” it under a set of identification assumptions. Athey et al. (2021) provide an example of how to generate simulation data to evaluate estimators of treatment effects under unconfoundedness, where they train two conditional WGANs to learn $X_{i}|D_{i}$ and $Y_{i}|(X_{i},D_{i})$ from the real sample. They then draw an artificial sample of treatment status and covariates from the WGAN-generated distribution of $X_{i}|D_{i}$, and for each $(X_{i},D_{i})$ drawn, an outcome is drawn from the WGAN-generated distribution of $Y_{i}|(X_{i},D_{i})$. Under unconfoundedness, $Y_{i}|(X_{i},D_{i}=d)$ identifies $Y_{i}(d)|X_{i}$ and hence for each $Y_{i}$ drawn, its counterfactual outcome is generated by mutating the corresponding treatment status $D_{i}$ to $(1-D_{i})$ but keeping the same $X_{i}$. In other words, the counterfactual outcome is drawn from the WGAN-generated distribution of $Y_{i}|(X_{i},1-D_{i})$. The true average treatment effect is simulated by averaging the difference between the pair of generated outcome and its counterfactual over a large WGAN-generated sample.
@@
Our setting, with a different design and identification assumption for the treatment effect of interest, requires a different way of using WGAN to generate suitable simulation data. Specifically, the treatment effects are only identified for $X_{i}\in\mathscr{B}$, so we need to simulate the treatment effect at each treatment boundary point of interest, instead of simply taking averages as in the unconfoundedness setting. To do so, we first train two conditional WGANs to learn $X_{i}|D_{i}$ and $Y_{i}|(X_{i},D_{i})$ from the real sample, as in Athey et al. (2021). But for a given treatment boundary point $x^{c}\in\mathscr{B}$, we draw a separate artificial sample to simulate the true treatment effect at $x^{c}$: after training the WGAN to learn the distribution of $Y_{i}|(X_{i},D_{i})$, we generate 100,000 outcomes from $Y_{i}|(X_{i}=x^{c},D_{i}=1)$ and another 100,000 outcomes from $Y_{i}|(X_{i}=x^{c},D_{i}=0)$, and take the average of their differences as the true treatment effect at $x^{c}$. [^8] We note that WGANs do not directly yield closed-form estimates of the fitted distribution, but rather generate a new, artificial sample of arbitrary size that closely mimics the real sample. This is why we need to simulate the true treatment effect at a given treatment boundary point $x^{c}\in\mathscr{B}$, unlike in the polynomial case where we can directly calculate the true treatment effects from the coefficients of the fitted polynomials evaluated at $x^{c}$. Figures 2 and 3 visualize the actual data and the simulated data from the polynomial specification and WGAN. In Appendix B, we include additional descriptions of the WGAN-generated DGPs in Figures B.1B.7, and we report in Table B.2 the Wasserstein distance between the real and WGAN data. We use the default options available from the wgan package of Athey et al. (2021) to train the WGANs; see Appendix B.3 for details. Following Section 5 of Athey et al. (2021), we conduct a thorough robustness check to assess the variability of the Monte Carlo results with respect to sampling variation and WGAN tuning parameters, with robustness results reported in Appendix B.5.
@@
Figure 2: Lee (2008): Democratic victory margin in the current election against the vote share in the next election. Panel A shows the observed data. Panel B plots the model data fitted using a 5-th order polynomial with different coefficients for the treated and control observations. Panel C plots the WGAN data. The cutoff is normalized to 0. The figure is generated using the rdrobust package with default smoothing (Calonico et al., 2014).
@@
Figure 3: Keele and Titiunik (2015): Geographic coordinates and election turnout. The leftmost panel shows the real data, and the middle and rightmost panels visualize respectively the simulated polynomial and WGAN DGPs. Coordinates below the black curve (treatment boundary) are treated, and the green circle shows the boundary point at which we estimate the treatment effect. To visualize the polynomial DGPs, 2,500 coordinates are sampled with replacement from the original sample plus a noise term drawn from $\mathcal{N}(0,0.01^{2})$.
@@ 4.2 Monte Carlo Results
We report the simulation results in Table 1. All simulations are implemented in R, where we use the rdrobust package of Calonico et al. (2014) for local linear regressions, the optrdd package of Imbens and Wager (2019) for the minimax-optimal estimator, and the grf package of Athey et al. (2019) for the two forest estimators—all with default options without tuning, except we specify pilot values for a few hyperparameters of the two forest estimators, for reasons explained in Appendix B.6.
@@
Table 1: Monte Carlo simulation results. For each method, we vary the sample size $n\in\{1000,5000,10000,50000,100000\}$ and report the bias, variance, and $95\%$ coverage rate across 1,000 Monte Carlo replications. Each panel corresponds to a different outcome variable in Section 4.1, with the true treatment effect in parentheses either calculated exactly (from the polynomial fit) or simulated as described in Section 4.1.1 (for the WGAN DGPs).
@@
Local linear regressions perform well in univariate settings, although they slightly undercover in the polynomial DGP (Panel A, Columns 1-5 of Table 1). [^9] This undercoverage is consistent with the simulation result ($90.6\%$ coverage) of Calonico et al. (2019), where they use the same polynomial DGP but with $n=1000$ for a total of $5000$ replications; see Table SA-1, Row 2 and Column 3 in their supplemental appendix. We have done our simulation using their STATA code and obtained similar results to those reported in Table 1. However, they are generally noisier in bivariate settings, where we follow Keele and Titiunik (2015) to first collapse the bivariate score using rescaled Euclidean (chordal) distance and then run rdrobust, especially in small samples. Although the variance of local linear regressions shrinks as the sample size gets large, their bias does not shrink fast enough relative to the diminishing variance in those bivariate WGAN DGPs (Panels F-H, Columns 1-5 of Table 1), and therefore the coverage rate drops as the sample size increases. This problem arises possibly due to the zero-density issue of the collapsed multivariate score at the new univariate cutoff, as we point out in Section 3.1.1 and empirically illustrate in the last two histograms of Figure 1, cautioning against such univariate transformations.
@@
The minimax-optimal estimator of Imbens and Wager (2019) has a relatively low bias but tends to overcover, which is expected since it is designed for minimizing the worst-case MSE and inference aims at achieving uniform coverage over a class of CEFs explained in Section 3.2. Nevertheless, its performance is contingent on how precisely the second derivative bound $\mathcal{B}$ is estimated: for the two artificial datasets built using the Lee (2008) data (Panels A-B of Table 1), the minimax-optimal estimator undercovers severely, especially for the polynomial DGP in Panel A. There, we follow the recommendation of Imbens and Wager (2019) to estimate the CEFs globally using a second-order polynomial and multiply the maximal estimated curvature by $4$. But this turns out to yield an estimate of $\mathcal{B}$ that is too small; Table B.1 in Appendix B.2 shows how the simulation results change with the constant multiplied by $\mathcal{B}$, and it is only after multiplying the maximal estimated curvature by $30$ do we get an estimate that is close to the true (least upper) bound $\mathcal{B}$.
Of course, one could argue that estimating $\mathcal{B}$ by a global second-order polynomial is a bad choice to start with. For the bivariate design, we follow their working paper (Imbens and Wager, 2018, p.21) and estimate $\mathcal{B}$ using the $95$-th percentile (over all sample points) of the operator norm curvature via a cross-validated ridge regression with interacted $7$-th order natural splines, which yields good coverage (Panels C-H, Columns 6-10 of Table 1). However, without knowledge about the true DGP, choosing a “good” model is not straightforward, and estimating second derivatives is more difficult when the dimension of the running variable increases. In addition, the current implementation of optrdd accepts up to two scores only, limiting its application more broadly.
@@
Honest regression forests and local linear forests have comparable performance to local linear regressions in the univariate design (Panels A-B, Columns 11-20 of Table 1), though their variances are slightly higher. On the other hand, regression forests are generally more biased than local linear forests, especially in small samples, which is a known boundary problem for regression forests. In addition, the variance of regression forests tends to be smaller than that of local linear forests, exacerbating the undercoverage problem. As we discussed in Section 3.3.2, local linear forests are better at capturing smooth signals via a linear correction, and its advantage over regression forests is especially pronounced when the underlying DGP exhibits strong linear trends near the boundary point at which we estimate the treatment effect. In Appendix B.5, we examine the robustness of the results in Table 1 to the number of training epochs, which is a tuning parameter of WGAN, and increasing the number of training epochs leads to stronger linear trends near the boundary point (see Figure B.7); in these cases, the performance of regression forests deteriorates, whereas local linear regressions have robust performance. With that said, these two forest-based methods are theoretically valid alternatives to existing methods for multivariate RD designs and can flexibly model any given number of scores, and empirical researchers may leverage and tailor them accordingly to fit different empirical settings.
5 Empirical Application
We revisit the empirical analysis in Narita and Yata (2023), who study the effect of hospital relief funding during COVID-19.
5.1 Data
The pandemic inflicted substantial financial pressure on U.S. hospitals, compounding revenue losses with increasing operational expenses. To alleviate this, the U.S. government enacted the Coronavirus Aid, Relief, and Economic Security (CARES) Act in March 2020 that provides emergency financial support to healthcare providers. A \$10 billion portion of this funding was allocated to “safety net” hospitals, as determined by a set of eligibility criteria that target hospitals serving vulnerable populations. Hospitals were eligible for safety net funding if they met all of the following three conditions: [^10] Source: Provider Relief Fund Frequently Asked Questions from the Health Resources and Services Administration, U.S. Department of Health and Human Services.
  1. 1.
    Medicare Disproportionate Patient Percentage (DPP) of 20.2% or higher. [^11] For a more detailed definition, see Centers for Medicare & Medicaid Services.
  2. 2.
    Annual Uncompensated Care (UCC) exceeding \$25,000 per bed. [^12] For a more detailed definition, see Fact Sheet: Uncompensated Hospital Care Cost.
  3. 3.
    Profit margin (defined as net income divided by the sum of net patient revenue and total other income) of 3.0% or less.
To determine safety net funding eligibility, we use data from the Healthcare Cost Report Information System (HCRIS) for the 2019 financial year (White, 2018). The 2019 measurements are the closest available approximation of the values hospitals likely reported when they applied for funding in 2020.131313Although using the 2020 data may seem more temporally aligned with the application timeline, it is unclear whether these records were actually used by the hospitals at the time of application. Since the coverage of the 2019 data was low at the time of Narita and Yata (2023)’s study, they use the 2018 HCRIS data instead. We thank them for their suggestion to use the updated data and help with data preparation. For the outcome variable, following Narita and Yata (2023), we use the number of confirmed or suspected COVID patients hospitalized in each hospital during the week spanning August 2 to August 8, 2020, drawn from the most recent (June 2024) version of the COVID Reported Patient Impact and Hospital Capacity by Facility dataset (U.S. Department of Health and Human Services, 2024). We note that, due to changes in the reporting format, our weekly outcome is based on a slightly different weakly reporting window from that of Narita and Yata (2023), which starts from July 31 to August 6, 2020 using an earlier version of the outcome dataset.
We follow the steps in Appendix H of Narita and Yata (2023) for data cleaning and missing value imputation. [^14] In addition, we remove one hospital that appears twice in the outcome data, as the two occurrences have considerably different values for the outcome variable. The resulting treated and control sample sizes are $n_{1}=656$ and $n_{0}=3$,199, respectively. Figure 4 visualizes the sample data as a scatter plot, where red (black) colored dots correspond to treated (control) hospitals; the treatment boundary is consisted of three partially bounded planes colored in pink. Under the assumptions in Section 2, we can identify the conditional treatment effect of funding eligibility at any point on the pink surface.
Figure 4: Visualization of the CARES Act data. The treatment boundary is shown in pink, with its edges colored in blue to represent variation along a single running variable while holding the other two at their respective cutoffs. The sample observations are displayed as a scatter plot, with dots colored in black (red) denoting control (treated) observations.
5.2 Empirical Results
Safety net funding eligibility criteria create a three-dimensional RD setup, where our framework applies. Acknowledging that directly estimating the conditional treatment effects nonparametrically at each treatment boundary point $X_{i}$ “is hard to use in practice, however, when $X_{i}$ has many elements (p. 12),” Narita and Yata (2023) propose a 2SLS estimator that is consistent for a weighted average of conditional treatment effects, nesting both propensity score matching and regression discontinuity in the case where the overlap condition fails in the first stage. In their empirical application, they use funding eligibility as an instrument for the amount of funding received and estimate the average effect of received funding amount on multiple outcomes, including the number of COVID patients hospitalized. Their findings suggest little to no effect. We find a similar null effect, but our analysis views funding eligibility as a binary treatment and directly estimates its causal effect. Moreover, we estimate conditional treatment effects as a function of treatment boundary point, allowing investigation of treatment effect heterogeneity along the treatment boundary. We find the estimated treatment effects, albeit with confidence intervals (CIs) covering zero, exhibit sizable heterogeneity along the treatment boundary.
We apply the honest regression forest and local linear forest estimators in (7) and (13) to the CARES Act data and, for ease of illustration, we plot how the treatment effect varies along each of the three running variables, while holding the other two fixed. This corresponds to estimating the treatment effect along the three “edges” of the treatment boundary, shown in blue in Figure 4. Due to the potential zero-density issue of local linear regressions with a collapsed univariate score, and the inability of the minimax-optimal estimator to accommodate more than two scores, we omit analyses using these two estimators.
Figure 5: Treatment effect estimates (blue lines) with $95$% CIs (blue shaded regions) for the honest regression forest estimator (top three panels) and the local linear forest estimator (bottom three panels).
For simplicity and better generalizability, we do not tune the hyperparameters of the forest estimators, but instead use the default values and set a few to specific pilot values as detailed in Section B.6. We use 50,000 trees for more stable results, as the grf package is known to be irreproducible across platforms. [^15] See the grf package manual. Figure 5 presents the estimation results with 95% CIs. Comparing across the top and bottom panels, we see little qualitative difference between the two estimators, and the CIs for the point estimates generally cover zero, supporting Narita and Yata (2023)’s conclusion that the relief funding was not well targeted toward hospitals most in need. There is, however, noticeable treatment effect heterogeneity as DPP and profit margin vary, suggesting that more vulnerable hospitals—those with higher DPP and lower profit margin—tend to utilize the relief funding more efficiently by providing care to more patients.